OpenStudy (anonymous):

Chris drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 6 hours. When Chris drove home, there was no traffic and the trip only took 4 hours. If his average rate was 22 miles per hour faster on the trip home, how far away does Chris live from the mountains?

OpenStudy (madhu.mukherjee.946):

Keisha drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 8 hours. When Keisha drove home, there was no traffic and the trip only took 6 hours. If her average rate was 16 miles per hour faster on the trip home, how far away does Keisha live from the mountains? way up distance from home = x speed = y time =8 hours d=rt x=8y x-8y=0...............(1) return distance =x speed = y+16 time = 6 hours x=6(y+16) x=6y+96 x-6y=96...............(2) x -8 y = 0 .............1

OpenStudy (madhu.mukherjee.946):

x -6 y = 96 .............2 Eliminate y multiply (1)by 6 Multiply (2) by 8 6 x -48 y = 0 8 x -48 y = 768 Add the two equations 14 x = 768 / 14 x = 55 plug value of x in (1) 1 x -8 y = 0 55 -8 y = 0 -8 y = 0 -55 -8 y = -55 y = 7 x= 55 the distance from home

OpenStudy (madhu.mukherjee.946):

this is an example of the sum you posted now pls try on ur own